A FABRY-PERÓT CAVITY PULSED FOURIER TRANSFORM W-BAND SPECTROMETER WITH A PULSED NOZZLE SOURCE

GARRY S. GRUBBS II, CHRISTOPHER T. DEWBERRY and \underline{STEPHEN A. COOKE}, Department of Chemistry, The University of North Texas, P. O. Box 305070, Denton, Texas, 76203, U.S.A.

We will present the design and operation of a simple Balle-Flygare-type spectrometer operable between 75 and 110 GHz (W-band). W-band frequencies are generated through the use of active multiplier chains. The circuit uses single side-band modulation with an IF of 180 MHz. The Fabry-Perót cavity consists of two opposing spherically concave mirrors, 70 mm in diameter with a 250 mm radius of curvature. The mirrors have been machined from copper and diamond polished to achieve a maximum surface defect of 135 nm. A solenoid valve admits pulses of gas into the cavity which is housed in a small vacuum chamber maintained at a resting pressure of <10^{−7} bar using a turbomolecular pump. The instruments use and sensitivity will be demonstrated using some simple gas phase species.